# Bootstrapping Vision-language Models for Self-supervised Remote Physiological Measurement

### v1

@@ Abstract
Facial video-based remote physiological measurement is a promising research area for detecting human vital signs (e.g. , heart rate, respiration frequency) in a non-contact way. Conventional approaches are mostly supervised learning, requiring extensive collections of facial videos and synchronously recorded photoplethysmography (PPG) signals. To tackle it, self-supervised learning has recently gained attentions; due to the lack of ground truth PPG signals, its performance is however limited. In this paper, we propose a novel self-supervised framework that successfully integrates the popular vision-language models (VLMs) into the remote physiological measurement task. Given a facial video, we first augment its positive and negative video samples with varying rPPG signal frequencies. Next, we introduce a frequency-oriented vision-text pair generation method by carefully creating contrastive spatio-temporal maps from positive and negative samples and designing proper text prompts to describe their relative ratios of signal frequencies. A pre-trained VLM is employed to extract features for these formed vision-text pairs and estimate rPPG signals thereafter. We develop a series of generative and contrastive learning mechanisms to optimize the VLM, including the text-guided visual map reconstruction task, the vision-text contrastive learning task, and the frequency contrastive and ranking task. Overall, our method for the first time adapts VLMs to digest and align the frequency-related knowledge in vision and text modalities. Extensive experiments on four benchmark datasets demonstrate that it significantly outperforms state of the art self-supervised methods.
Index Terms:
Remote physiological measurement, vision-language models, contrastive learning, generative learning.
@@ 1 Introduction
Recently, vision-language models (VLMs), such as contrastive language–image pre-training (CLIP) [25], video and language understanding (VindLU) [26], have drawn increasing research attentions due to their remarkable ability in many vision and language tasks. Particularly, in video processing, given vision-text pairs, visual and textual embeddings are extracted by vision and text encoders, respectively; and are then aligned via vision-text contrastive learning. As shown in Fig. 1, the text prompts can be versatile, i.e. ”the color of sky changes from gold to orange to blue”, covering rich temporal details. This type of text cues thus facilitates the model understanding of temporal visual patterns in the videos. Therefore, such VLMs have showed special strength in temporal tasks like action recognition [27], video summarization [28] and temporal action localization [29]. Inspired by their successes, we aim to leverage the powerful temporal feature modeling ability of pre-trained VLMs to capture the periodic variation of skin color in facial videos. As far as we have explored, VLMs have not been used to capture the frequency-related attributes in images/videos so far. Intuitively, we posit two potential solutions, specified below.
@@
In this paper, we introduce a pioneering approach to bootstrap VLMs for self-supervised remote physiological measurement. Our proposed method, namely VL-phys, marks the first attempt to adapt VLMs with the ability to digest the frequency-related knowledge in vision and text modalities. As shown in Fig. 1, we introduce a novel frequency-oriented generation mechanism to create vision-text pairs reflecting the frequencies of skin color temporal variations. Given the input facial video, we first augment it into multiple positive and negative samples with their rPPG frequencies conforming certain relations; next, in the absence of their ground truth signal frequencies, we create contrastive spatio-temporal maps between positive and negative samples and craft text prompts to describe the frequency relations across samples; this forms vision-text pairs. Afterwards, we fine-tune the pre-trained VLM with these created vision-text pairs through well-designed representation learning in both generative and contrastive manners. The former features a text-guided visual map reconstruction task in which we reconstruct masked patches of contrastive spatio-temporal maps under the guidance of text prompts; while the latter consists of a series of tasks, such as the multimodal vision-text contrastive learning task to align the embeddings of vision-text pairs, and the unimodal frequency contrastive and ranking task to optimize rPPG signals estimated from visual embeddings of different video samples.
@@
  • We propose VL-phys, a novel self-supervised framework for remote physiological measurement based on the pre-trained VLM. It innovatively adapt the pre-trained VLM to digest the frequency-related knowledge in vision and text modalities via representation learning in both contrastive and generative manners.
@@
  • We design a multitask learning pipeline that incorporates the multimodal generative and contrastive tasks (i.e. , the text-guided visual map reconstruction task and the vision-text contrastive learning task) to help optimize the self-supervised rPPG estimation. Besides the multimodal learning, we also introduce the unimodal frequency contrastive and ranking task to constrain rPPG estimation among multiple augmented video samples.
@@ 2.1 Remote physiological measurement
Recently, deep learning approaches show remarkable performance in rPPG estimation [6, 5, 15, 16, 17, 19, 18]. For example, Niu et al. [19] proposed a feature disentangling mechanism that effectively separates physiological features from noises by employing an encoder-decoder network. Yu et al. [6] utilized self-attention layers to capture long-range spatio-temporal relationships among multiple video clips for rPPG estimation. Gupta et al. [38] based their work on vision transformer to regress rPPG waveforms. Despite their remarkable achievements, these supervised learning approaches demand extensive facial videos and synchronously recorded PPG signals for model training. The collection of PPG signals is both costly and time-consuming. Consequently, there has been a recent surge in interest to optimize the model using contrastive self-supervised learning [20, 21, 22, 23]. Specifically, they first apply spatial/frequency augmentations to create the positive/negative samples for the given video. Next, they pull close the predicted rPPG signals from positive samples, and repel the predictions from positive samples to negative samples. For instance, Yue et al. [21] designed a frequency augmentation module to modify the frequency of the rPPG signal for the given video and applied contrastive learning between the transformed and original videos. Similarly, Sun et al. [23] presented Contrast-Phys+ which contrasts rPPG similarity among different spatial crops and temporal clips within the given video. Due to the absence of supervision from PPG signals, learning the periodic temporal information is still very challenging for the self-supervised learning.
@@
Figure 2: Overall architecture of our VL-phys. Given an input video, we respectively apply spatial augmentation and learnable frequency augmentation (LFA) to obtain its positive and negative video samples. We generate their spatio-temporal maps (STMaps) and then create contrastive spatio-temporal maps (C-STMaps) to reflect the frequency ratios of skin color temporal variations between positive and negative samples; meanwhile, we carefully craft text prompts to describe such relations. Afterwards, we fine-tune the pre-trained vision and text encoders of VLM with these formed vision-text pairs via multimodal generative and contrastive tasks, i.e. the text-guided visual map reconstruction task and vision-text contrastive learning task. Moreover, we introduce the unimodal frequency contrastive loss and the frequency ranking loss to optimize the rPPG signals estimated from different video samples.
@@ 3.1 Overview
The method overview is illustrated in Fig. 2. Our goal is to bootstrap and finetune VLM to learn rPPG signals from unlabelled facial videos using generative and contrastive learning mechanisms. The framework contains five phases: data augmentation, vision-text pair generation, feature extraction, visual map reconstruction and network optimization.
Specifically, given a facial video $x^{a}$, we first perform spatial/frequency augmentation on it to create a number of positive/negative samples. The positive samples $X^{p}=\left\{x^{p}\right\}$ maintain the same rPPG signal frequency ($f^{a}$) to $x^{a}$, while the rPPG signal frequencies of negative samples are transformed to $r\times f^{a}$ by feeding a set of frequency ratios $R=\left\{r\right\}$ into a learnable frequency augmentation (LFA) module. Then we generate the spatio-temporal maps (STMaps) $M^{p}=\left\{m^{p}\right\}$ and $M^{n}=\left\{m^{n}\right\}$ for both positive and negative samples. These STMaps represent the temporal variations of skin colors in facial ROIs. Next, we introduce a novel frequency-oriented vision-text pair generation method: we create a set of contrastive spatio-temporal maps (C-STMaps) $M^{c}=\left\{m^{c}\right\}$ by concatenating any $m^{p}$ with $m^{n}$ horizontally. Since the frequency ratio of rPPG signals between $m^{p}$ with $m^{n}$ is a fixed value $r$, we construct the text prompt $d\in D$ that depicts the relative frequency ratio of the color variation from left to right in $m^{c}$. By this means, $m^{c}$ and $d$ form a vision-text pair. Meanwhile, several non-overlapping patches of $m^{c}$ are randomly masked to generate the masked contrastive spatio-temporal maps (M-STMap) $m^{o}\in M^{o}$. Having $M^{p}$, $M^{n}$, $M^{c}$ and $M^{o}$, in the third stage, we pass them into the vision encoder to obtain visual embeddings $V^{p}$, $V^{n}$, $V^{c}$ and $V^{o}$. Meanwhile, the text prompts in $D$ are input into the text encoder to obtain textual embeddings. We leverage the pre-trained vision and text encoders from the popular VLM, VindLU [26]. After that, an rPPG estimation head is applied to project $V^{p}$, $V^{n}$ into rPPG signals $Y^{p}=\left\{y^{p}\right\}$ and $Y^{n}=\left\{y^{n}\right\}$. In the visual map reconstruction stage, we propose a text-guided visual map reconstruction (TVR) module to reconstruct the masked patches in $m^{o}$ under the guidance of text prompts. Last, in the main network optimization stage, the vision-text contrastive loss [25] is applied to align the embeddings of every vision-text pair $m^{c}$ and $d$. The frequency contrastive loss [21] is exploited among the power spectral densities (PSD) of signals $Y^{p}$ and $Y^{n}$. Also, we introduce a novel frequency ranking loss among $Y^{p}$ and $Y^{n}$ to let the vision encoder accurately predict their frequency ranks.
@@ 3.3.2 Frequency-oriented vision-text pair generation
Next, we generate text prompt $d\in D$ to respectively describe the relative ratio of signal frequencies between the left and right sides of each $m^{c}$. Specifically, $d$ is defined as: “the frequency of the horizontal color variation on the left/right side is [ratio] times of that on the right/left side of the image”. Here, the [ratio] token is replaced with the specific frequency ratio $r$ between negative and positive samples. For the example shown in Fig. 3, the [ratio] token is replaced with the value “1/2”. By this means, the C-STMap $m^{c}_{j}$ and its corresponding text prompt $d_{j}$ effectively form a positive vision-text pair. In contrast, $m^{c}_{j}$ combined with other C-STMaps’ text prompts $\{d_{l}|l\in\left[1,6\right],l\neq j\}$ form negative vision-text pairs. These pairs are used for the subsequent vision-text contrastive learning and text-guided visual map reconstruction. The experimental analysis of other text templates is presented in Sec. 6.3.
@@ 3.4 Feature extraction
Figure 4: The structure of text-guided visual map reconstruction (TVR) module.
3.5 Visual map reconstruction
In this stage, we enable the encoders to develop a deeper understanding of the frequency-related attributes via a generative learning mechanism. We develop a text-guided visual map reconstruction task to reconstruct masked patches of C-STMaps with the guidance of the text prompts. This task can implicitly enhance our vision encoder to learn the periodicity of rPPG signals. Below we specify the details.
@@ 3.5.1 Masked contrastive spatio-temporal map generation
3.5.2 Text-guided visual map reconstruction
We further design a text-guided visual map reconstruction (TVR) module to recover the masked patches in $m^{o}$. Specifically, the TVR module contains three interaction blocks. Its structure is illustrated in Fig. 4. The input for the TVR module, $z^{o}$, is the concatenated sequence of encoded visible patch features $v^{o}$ and the learnable masked tokens (with positional embeddings). Each interaction block begins by calculating the self-attention (SA) within $z^{o}$. It is expressed as: $v_{sa}^{o}=SA(LN(z^{o}))+z^{o}$, where $LN(.)$ represents the layer normalization. Then we apply the cross-attention (CA) between $v_{sa}^{o}$ and its associated textual embedding $\hat{d}$ to facilitate their interaction and obtain joint representation $v_{ca}^{o}$; $v_{sa}^{o}$ is treated as the query and $\hat{d}$ is treated as both the key and value for CA. This process is written as: $v_{ca}^{o}=CA(LN(v_{sa}^{o}),\hat{d})+v_{sa}^{o}$. Subsequently, a feed-forward layer is exploited to further refine $v_{ca}^{o}$. After executing three such blocks, we reconstruct the patches and map the output back to the original RGB space via a $1\times 1$ convolution layer and an upsampling layer.
@@ 3.6 Network optimization
Vision-text contrastive loss ($L_{vtc}$): This loss aims to align text with vision in the embedding space. This is achieved by encouraging the similarities between visual and textual embeddings from positive vision-text pairs (i.e. $m^{c}_{j}$ and $d_{j}$) while distinguishing them from negative pairs (i.e. $m^{c}_{j}$ and {$d_{l}\}$) (see Sec. 3.3.2). $L_{vtc}$ has been demonstrated to be an effective objective for vision-text contrastive learning. We define it as:
@@
TABLE I: Comparison to state of the art on HR estimation. The results are reported on UBFC-rPPG, PURE, VIPL-HR, and MMVS datasets. $\uparrow$ indicates that the larger the value is the better it is and $\downarrow$ vice versa. The best supervised approach is marked in shadow, while the best self-supervised approach is marked in bold.
Method PPG UBFC-rPPG PURE VIPL-HR MMVS
[…]
POS [12] - 8.35 10.00 0.24 3.14 10.57 0.95 11.53 17.26 0.30 6.77 9.40 0.82
CHROM [14] - 8.20 9.92 0.27 3.82 6.80 0.97 11.42 16.92 0.28 6.85 9.37 0.82
Green [3] - 6.01 7.87 0.29 4.39 11.60 0.90 - - - 7.13 9.46 0.80
SynRhythm [48] 5.59 6.82 0.72 2.71 4.86 0.98 - - - 4.48 6.52 0.89
Meta-rppg [49] 5.97 7.42 0.53 2.52 4.63 0.98 - - - 4.30 6.20 0.91
PulseGan [50] 1.19 2.10 0.98 2.28 4.29 0.99 - - - 3.52 5.09 0.93
Dual-Gan [5] 0.44 0.67 0.99 0.82 1.31 0.99 4.93 7.68 0.81 3.00 4.27 0.94
Physformer [6] 0.40 0.71 0.99 1.10 1.75 0.99 4.97 7.79 0.78 3.28 4.50 0.93
Du et al. [7] 0.16 0.57 0.99 - - - - - - - - -
Dual-TL [51] 0.17 0.41 0.99 0.37 0.68 0.99 6.38 6.92 0.69 - - -
ND-DeeprPPG [52] 0.31 0.98 0.99 0.18 0.41 0.99 - - - - - -
Gideon et al. [22] 1.85 4.28 0.93 2.32 2.97 0.99 9.80 15.48 0.38 3.43 4.74 0.93
Yue et al. [21] 0.58 0.94 0.99 1.23 2.01 0.99 - - - 2.93 4.16 0.94
Contrast-Phys+ [23] 0.64 1.00 0.99 1.00 1.40 0.99 7.49 14.40 0.49 2.96 4.11 0.94
SiNC [53] 0.59 1.83 0.99 0.61 1.84 0.99 - - - 3.15 4.40 0.93
VL-phys 0.28 0.60 0.99 0.52 0.95 0.99 6.04 8.78 0.71 2.33 3.76 0.95

### v2

@@ Abstract
Facial video-based remote physiological measurement is a promising research area for detecting human vital signs (e.g. , heart rate, respiration frequency) in a non-contact way. Conventional approaches are mostly supervised learning, requiring extensive collections of facial videos and synchronously recorded photoplethysmography (PPG) signals. To tackle it, self-supervised learning has recently gained attentions; due to the lack of ground truth PPG signals, its performance is however limited. In this paper, we propose a novel frequency-centric self-supervised framework that successfully integrates the popular vision-language models (VLMs) into the remote physiological measurement task. Given a facial video, we first augment its positive and negative video samples with varying rPPG signal frequencies. Next, we introduce a frequency-oriented vision-text pair generation method by carefully creating contrastive spatio-temporal maps from positive and negative samples and designing proper text prompts to describe their relative ratios of signal frequencies. A pre-trained VLM is employed to extract features for these formed vision-text pairs and estimate rPPG signals thereafter. We develop a series of frequency-related generative and contrastive learning mechanisms to optimize the VLM, including the text-guided visual reconstruction task, the vision-text contrastive learning task, and the frequency contrastive and ranking task. Overall, our method for the first time adapts VLMs to digest and align the frequency-related knowledge in vision and text modalities. Extensive experiments on four benchmark datasets demonstrate that it significantly outperforms state of the art self-supervised methods.
keywords:
Remote physiological measurement, vision-language models, frequency-related generative and contrastive learning, facial video analysis.
@@ 1 Introduction
Recently, vision-language models (VLMs), such as contrastive language–image pre-training (CLIP) [42], video and language understanding (VindLU) [6], have drawn increasing research attentions due to their remarkable ability in many vision and language tasks [76, 72, 44]. Particularly, in video processing, given vision-text pairs, visual and textual embeddings are extracted by vision and text encoders, respectively; and are then aligned via vision-text contrastive learning. As shown in Fig. 1, the text prompts can be versatile, i.e. ”the color of sky changes from gold to orange to blue”, covering rich temporal details. This type of text cues thus facilitates the model understanding of temporal visual patterns in the videos. Therefore, such VLMs have showed special strength in temporal tasks like action recognition [55], video summarization [39] and video object segmentation [68]. Inspired by their successes, we aim to leverage the powerful temporal feature modeling ability of pre-trained VLMs to capture the periodic variation of skin color in facial videos. As far as we have explored, VLMs have not been used to capture the frequency-related attributes in images/videos so far. Intuitively, we posit two potential solutions, specified below.
@@
In this paper, we introduce a pioneering approach to bootstrap VLMs for self-supervised remote physiological measurement. Our proposed method, namely VL-phys, marks the first attempt to adapt VLMs with the ability to digest the frequency-related knowledge in vision and text modalities. As shown in Fig. 1, we introduce a novel frequency-oriented generation mechanism to create vision-text pairs reflecting the frequencies of skin color temporal variations. Given the input facial video, we first augment it into multiple positive and negative samples with their rPPG frequencies conforming certain relations; next, in the absence of their ground truth signal frequencies, we create contrastive spatio-temporal maps between positive and negative samples and craft text prompts to describe the frequency relations across samples; this forms vision-text pairs. Afterwards, we fine-tune the pre-trained VLM with these created vision-text pairs through well-designed frequency-related representation learning in both generative and contrastive manners. The former features a text-guided visual reconstruction task in which we reconstruct masked patches of contrastive spatio-temporal maps under the guidance of text prompts; while the latter consists of a series of tasks, such as the multimodal vision-text contrastive learning task to align the embeddings of vision-text pairs, and the unimodal frequency contrastive and ranking task to optimize rPPG signals estimated from visual embeddings of different video samples.
@@
  • We propose VL-phys, a novel frequency-centric self-supervised framework for remote physiological measurement based on the pre-trained VLM. It innovatively adapt the pre-trained VLM to digest the frequency-related knowledge in vision and text modalities via representation learning in both contrastive and generative manners.
@@
  • We design a multitask learning pipeline that incorporates the multimodal generative and contrastive tasks (i.e. , the text-guided visual reconstruction task and the vision-text contrastive learning task) to help optimize the self-supervised rPPG estimation. Besides the multimodal learning, we also introduce the unimodal frequency contrastive and ranking task to constrain rPPG estimation among multiple augmented video samples.
@@
Our method, beyond its merits in the remote physiological measurement task, should also benefit many other tasks that desire VLMs to understand the periodic and frequency-related attributes in their modalities, such as gait recognition [30], weather change prediction [45], Parkinson’s disease diagnosis [61].
@@ 2.1 Remote physiological measurement
Recently, deep learning approaches show remarkable performance in rPPG estimation [67, 32, 41, 19, 65, 37, 24, 53, 25]. For example, Niu et al. [37] proposed a feature disentangling mechanism that effectively separates physiological features from noises by employing an encoder-decoder network. Yu et al. [67] utilized self-attention layers to capture long-range spatio-temporal relationships among multiple video clips for rPPG estimation. Gupta et al. [12] based their work on vision transformer to regress rPPG waveforms. Despite their remarkable achievements, these supervised learning approaches demand extensive facial videos and synchronously recorded PPG signals for model training. The collection of PPG signals is both costly and time-consuming. Consequently, there has been a recent surge in interest to optimize the model using contrastive self-supervised learning [62, 71, 10, 49]. Specifically, they first apply spatial/frequency augmentations to create the positive/negative samples for the given video. Next, they pull close the predicted rPPG signals from positive samples, and repel the predictions from positive samples to negative samples. For instance, Yue et al. [71] designed a frequency augmentation module to modify the frequency of the rPPG signal for the given video and applied contrastive learning between the transformed and original videos. Similarly, Sun et al. [49] presented Contrast-Phys+ which contrasts rPPG similarity among different spatial crops and temporal clips within the given video. Due to the absence of supervision from PPG signals, learning the periodic temporal information is still very challenging for the self-supervised learning.
@@
Figure 2: Overall architecture of our VL-phys. Given an input video, we respectively apply spatial augmentation and learnable frequency augmentation (LFA) to obtain its positive and negative video samples. We generate their spatio-temporal maps (STMaps) and then create contrastive spatio-temporal maps (C-STMaps) to reflect the frequency ratios of skin color temporal variations between positive and negative samples; meanwhile, we carefully craft text prompts to describe such relations. Afterwards, we fine-tune the pre-trained vision and text encoders of VLM with these formed vision-text pairs via frequency-related multimodal generative and contrastive tasks, i.e. the text-guided visual reconstruction task and vision-text contrastive learning task. Moreover, we introduce the unimodal frequency contrastive loss and the frequency ranking loss to optimize the rPPG signals estimated from different video samples.
@@ 3.1 Overview
The method overview is illustrated in Fig. 2. Our goal is to bootstrap and finetune VLM to learn rPPG signals from unlabelled facial videos using frequency-related generative and contrastive learning mechanisms. The framework contains five phases: data augmentation, vision-text pair generation, feature extraction, visual reconstruction and network optimization.
Specifically, given a facial video $x^{a}$, we first perform spatial/frequency augmentation on it to create a number of positive/negative samples. The positive samples $X^{p}=\left\{x^{p}\right\}$ maintain the same rPPG signal frequency ($f^{a}$) to $x^{a}$, while the rPPG signal frequencies of negative samples are transformed to $r\times f^{a}$ by feeding a set of frequency ratios $R=\left\{r\right\}$ into a learnable frequency augmentation (LFA) module. Then we generate the spatio-temporal maps (STMaps) $M^{p}=\left\{m^{p}\right\}$ and $M^{n}=\left\{m^{n}\right\}$ for both positive and negative samples. These STMaps represent the temporal variations of skin colors in facial ROIs. Next, we introduce a novel frequency-oriented vision-text pair generation method: we create a set of contrastive spatio-temporal maps (C-STMaps) $M^{c}=\left\{m^{c}\right\}$ by concatenating any $m^{p}$ with $m^{n}$ horizontally. Since the frequency ratio of rPPG signals between $m^{p}$ with $m^{n}$ is a fixed value $r$, we construct the text prompt $d\in D$ that depicts the relative frequency ratio of the color variation from left to right in $m^{c}$. By this means, $m^{c}$ and $d$ form a vision-text pair. Meanwhile, several non-overlapping patches of $m^{c}$ are randomly masked to generate the masked contrastive spatio-temporal maps (M-STMap) $m^{o}\in M^{o}$. Having $M^{p}$, $M^{n}$, $M^{c}$ and $M^{o}$, in the third stage, we pass them into the vision encoder to obtain visual embeddings $V^{p}$, $V^{n}$, $V^{c}$ and $V^{o}$. Meanwhile, the text prompts in $D$ are input into the text encoder to obtain textual embeddings. We leverage the pre-trained vision and text encoders from the popular VLM, VindLU [6]. After that, an rPPG estimation head is applied to project $V^{p}$, $V^{n}$ into rPPG signals $Y^{p}=\left\{y^{p}\right\}$ and $Y^{n}=\left\{y^{n}\right\}$. In the visual reconstruction stage, we propose a text-guided visual reconstruction (TVR) module to reconstruct the masked patches in $m^{o}$ under the guidance of text prompts. Last, in the main network optimization stage, the vision-text contrastive loss [42] is applied to align the embeddings of every vision-text pair $m^{c}$ and $d$. The frequency contrastive loss [71] is exploited among the power spectral densities (PSD) of signals $Y^{p}$ and $Y^{n}$. Also, we introduce a novel frequency ranking loss among $Y^{p}$ and $Y^{n}$ to let the vision encoder accurately predict their frequency ranks.
@@ 3.3.1 Contrastive spatio-temporal map generation
Note that here we use the horizontal concatenation to ensure that each row corresponding to the same ROI in the positive and negative STMaps are properly aligned. This process yields the C-STMap in which the frequency of horizontal color variation displays a noticeable change at the boundary of the two sides. This arrangement allows us to easily design text prompts to describe the relative ratio of signal frequencies between the left and right sides of C-STMap. In contrast, using other concatenation methods, such as vertical or channel concatenation, would alter the spatial alignment between the positive and negative STMaps, and the signal frequency difference would not be manifesting to the same extent as along the single horizontal axis of C-STMap. This would make it even undesirable to generate corresponding text prompts for the frequency-related representation learning.
@@ 3.3.2 Frequency-oriented vision-text pair generation
Next, we generate text prompt $d\in D$ to respectively describe the relative ratio of signal frequencies between the left and right sides of each $m^{c}$. Specifically, $d$ is defined as: “the frequency of the horizontal color variation on the left/right side is [ratio] times of that on the right/left side of the image”. Here, the [ratio] token is replaced with the specific frequency ratio $r$ between negative and positive samples. For the example shown in Fig. 3, the [ratio] token is replaced with the value “1/2”. By this means, the C-STMap $m^{c}_{j}$ and its corresponding text prompt $d_{j}$ effectively form a positive vision-text pair. In contrast, $m^{c}_{j}$ combined with other C-STMaps’ text prompts $\{d_{l}|l\in\left[1,6\right],l\neq j\}$ form negative vision-text pairs. These pairs are used for the subsequent vision-text contrastive learning and text-guided visual reconstruction. The experimental analysis of other text templates is presented in Sec. 6.4.
@@ 3.4 Feature extraction
Figure 4: The structure of text-guided visual reconstruction (TVR) module.
3.5 Visual reconstruction
In this stage, we enable the encoders to develop a deeper understanding of the frequency-related attributes via a frequency-related generative learning mechanism. We develop a text-guided visual reconstruction task to reconstruct masked patches of C-STMaps with the guidance of the text prompts. This task can implicitly enhance our vision encoder to learn the periodicity of rPPG signals. Below we specify the details.
@@ 3.5.1 Masked contrastive spatio-temporal map generation
3.5.2 Text-guided visual reconstruction
We further design a text-guided visual reconstruction (TVR) module to recover the masked patches in $m^{o}$. Specifically, the TVR module contains three interaction blocks. Its structure is illustrated in Fig. 4. The input for the TVR module, $z^{o}$, is the concatenated sequence of encoded visible patch features $v^{o}$ and the learnable masked tokens (with positional embeddings). Each interaction block begins by calculating the self-attention (SA) within $z^{o}$. It is expressed as: $v_{sa}^{o}=SA(LN(z^{o}))+z^{o}$, where $LN(.)$ represents the layer normalization. Then we apply the cross-attention (CA) between $v_{sa}^{o}$ and its associated textual embedding $\hat{d}$ to facilitate their interaction and obtain joint representation $v_{ca}^{o}$; $v_{sa}^{o}$ is treated as the query and $\hat{d}$ is treated as both the key and value for CA. This process is written as: $v_{ca}^{o}=CA(LN(v_{sa}^{o}),\hat{d})+v_{sa}^{o}$. Subsequently, a feed-forward layer is exploited to further refine $v_{ca}^{o}$. After executing three such blocks, we reconstruct the patches and map the output back to the original RGB space via a $1\times 1$ convolution layer and an upsampling layer.
@@ 3.6 Network optimization
Vision-text contrastive loss ($L_{vtc}$): This loss aims to align text with vision in the embedding space. This is achieved by encouraging the similarities between visual and textual embeddings from positive vision-text pairs (i.e. $m^{c}_{j}$ and $d_{j}$) while distinguishing them from negative pairs (i.e. $m^{c}_{j}$ and {$d_{l}\}$) (see Sec. 3.3.2). We define it as: